Ienoring leap-years and assuming all birthdays are equally likely, what is the probability that a persons two grandfathers have birthdays that are exactly two days apart.

PLEASE WRITE LEGIBLE
Answer #1

here number of ways to select 2 birthdays from 365 days =3652

number of ways two birthdays are 2 days apart =365*2 (as first birthday can be on any of 365 days and for second birthday we have 2 choices one 2 days later then first person birthday and other 2 days after first person birthday)

hence P(birthdays are exactly two days apart)=(365*2)/(365)2 =2/365